OpenStudy (anonymous):

How would I integrate (1-2x)^3?

OpenStudy (freckles):

do you know how to integrate something like: \[-\frac{1}{2}u^3\] w.r.t to u ?

OpenStudy (freckles):

\[u=1-2x \\ \frac{du}{dx}=-2 \\ du=-2 dx \\ \frac{-1}{2} du=dx \\ \int\limits (1-2x)^3 dx=\int\limits u^3 \frac{-1}{2} du=\frac{-1}{2} \int\limits u^3 du\]

OpenStudy (anonymous):

I understand until the 4th step. Would you be able to explain it to me?

OpenStudy (freckles):

the 4th line I wrote? I divide both sides by -2 or the equivalent multiplied both sides by -1/2

OpenStudy (anonymous):

Okay. :) Then do we use the chain rule for the 5th line you wrote?

OpenStudy (freckles):

yeah

OpenStudy (anonymous):

Do you need to simplify -1/2 (1-2x)^3 to get the final answer?

OpenStudy (freckles):

well you still need to integrate u^3

OpenStudy (freckles):

w.r.t. u

OpenStudy (anonymous):

So the next step would be -1/2 x u^4/4? And then I substitute 1-2x for u?

OpenStudy (freckles):

yep

OpenStudy (freckles):

and that one x is a multiplying symbol right?

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

\[\frac{-1}{2} \cdot \frac{u^4}{4}+C\] and yes you just replace u with (1-2x)

OpenStudy (freckles):

and you could rewrite as \[\frac{-(1-2x)^4}{8}+C\] I would do nothing else expanding that makes it look nastier

OpenStudy (anonymous):

So the final answer is -((1-2x)^4))/8 + c?

OpenStudy (anonymous):

Thank you! :)

OpenStudy (freckles):

np